A $100 par value 10-year bond provides coupons at 5% convertible semiannually. The yield rate is 4%

convertible semiannually. What is the flat price 8.4 years after issue at the same yield rate (assuming

compound interest)?
Answer #1

Given Face value = $100, r 1.6*2 3.2 5/2 2.5%, YTM = 4/2 =296, n Value of bond-interest PVIFA(2%,3.2) Value* PVIF(2%, 3.2) +